Find the lateral area and volume of the frustum of a right circular cone having a slant height of 14m and the radii of the bases are 4m and 1.5 m respectively.


1
Expert's answer
2021-12-21T16:19:29-0500

The lateral area of the frustum of a right circular cone is equal to one-half of the sum of the circumferences of the bases multiplied by slant height


"Lateral \ area=\dfrac{1}{2}(2\pi R+2\pi r)l"

"=\pi(R+r)l=\pi(4+1.5)(14)(m^2)=77\pi\ m^2"

The frustum volume is


"V=\dfrac{1}{3}\pi h(R^2+r^2+Rr)"

"h=\sqrt{l^2-(r-r)^2}=\sqrt{14^2-(4-1.5)^2}(m)"

"=\dfrac{\sqrt{759}}{2}(m)"

"V=\dfrac{1}{3}\pi (\dfrac{\sqrt{759}}{2})(4^2+1.5^2+4(1.5))(m^3)"

"=\dfrac{97\sqrt{759}}{24}\pi (m^3)"